OpenStudy (anonymous):

y(y-3) -0 solve for y!

OpenStudy (anonymous):

it should be =0 not minus

OpenStudy (anonymous):

Is that all to the equation? Is there another side of the equation. y(y-3)-0=?

OpenStudy (anonymous):

Ah ok.

OpenStudy (anonymous):

Start by distributing the y.

OpenStudy (anonymous):

so y^2 - 3y=0

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@Dscdago are you still there?

OpenStudy (anonymous):

Yes. I'm trying different ways to solve this. And at the same time writing it in here.

OpenStudy (anonymous):

ohh okay thank you! I thought you left haha

OpenStudy (anonymous):

You can maybe apply the quadratic formula. You have y^2-3y+0=0. \[\frac{ -(-3)\pm \sqrt{(-3)^{2}-4(1)(0)} }{ 2(1) }\]\[\frac{ -(-3)\pm \sqrt{(9)-0}}{ 2 }\]\[\frac{ -(-3)\pm \sqrt{9} }{ 2 }\]\[\frac{ -(-3)\pm 3 }{ 2 }\]\[\frac{ -(-3) + 3 }{ 2 }=\frac{ 3+3 }{ 2 }= \frac{ 6 }{ 2 }=3\]\[\frac{ -(-3)-3 }{ 2 }=\frac{ 3-3 }{ 2 }=\frac{ 0 }{ 2 }=0\] The solutions for the problem is y=0,3. Sorry it took long, I had a problem in my math and had to correct it.

OpenStudy (anonymous):

Oh my gosh! Thank you so so much!!!!

OpenStudy (anonymous):

No problem. Glad I could help.